# BeamLoRA: Beam-Constraint Low-Rank Adaptation

### v1

@@ Abstract
Due to the demand for efficient fine-tuning of large language models, Low-Rank Adaptation (LoRA) has been widely adopted as one of the most effective parameter-efficient fine-tuning methods. Nevertheless, while LoRA improves efficiency, there remains room for improvement in accuracy. Herein, we adopt a novel perspective to assess the characteristics of LoRA ranks. The results reveal that different ranks within the LoRA modules not only exhibit varying levels of importance but also evolve dynamically throughout the fine-tuning process, which may limit the performance of LoRA. Based on these findings, we propose BeamLoRA, which conceptualizes each LoRA module as a beam where each rank naturally corresponds to a potential sub-solution, and the fine-tuning process becomes a search for the optimal sub-solution combination. BeamLoRA dynamically eliminates underperforming sub-solutions while expanding the parameter space for promising ones, enhancing performance with a fixed rank. Extensive experiments across three base models and 12 datasets spanning math reasoning, code generation, and commonsense reasoning demonstrate that BeamLoRA consistently enhances the performance of LoRA, surpassing the other baseline methods. We will release our code soon.
@@ 1 Introduction
In recent years, large language models have shown tremendous potential in various applications (Touvron et al., 2023a, b; Jiang et al., 2023; OpenAI, 2023). To further enhance model performance on specific downstream tasks, fine-tuning is usually the most effective approach. However, as the scale of models keeps increasing, fine-tuning all model parameters becomes unsustainable. To address this issue, parameter-efficient fine-tuning (PEFT) emerges as a practical solution (Houlsby et al., 2019; Li and Liang, 2021; Liu et al., 2022; Hu et al., 2022). By updating only lightweight modules, these methods nearly achieve the results of full parameter fine-tuning while reducing both fine-tuning time and memory usage.
@@ 2.1 Low-Rank Adaptation (LoRA)
where $x\in{\mathbb{R}^{d}}$ represents the input and $y\in{\mathbb{R}^{d}}$ is the output. During fine-tuning, $\mathbf{W}_{0}$ remains frozen, while only $\mathbf{B}$ and $\mathbf{A}$ matrices are trainable.
@@ 2.2 Analysis of LoRA Ranks
Spatial Dimension. Figure 1 describes the results of sorting the importance of ranks after fine-tuning. The blue line represents the levels of importance across different deciles after sorted by importance in a LoRA module. It can be observed that the deciles of the importance of ranks are hierarchical, indicating that the importance of different ranks within a LoRA module is not uniform after fine-tuning. Furthermore, by pruning the ranks in each LoRA module based on its own importance from the least to the most significant gradually (the red line), the accuracy shows limited change when pruning the less important ranks. On the contrary, when important ranks are pruned, the evaluation results drop sharply to zero. This phenomenon further demonstrates the significant differences in importance among the different ranks.
@@
Figure 1: Differences in importance among ranks within a LoRA module (spatial). The blue line represents the deciles of importance for ranks. The red line represents accuracy changes when pruning ranks of varying importance. We take ffn.up_proj in the 30th layer as an example, with similar phenomena in other modules.
@@
Figure 2: The visualization of importance difference among different ranks in LoRA on LLaMA2-7B as fine-tuning steps increase (temporal). We take ffn.up_proj and attn.v_proj in the 30th layer as examples, with similar trends observed in all other modules.
@@
In summary, there are significant differences in the importance of ranks in LoRA, and these differences appear and gradually increase as the fine-tuning process progresses. However, in most existing LoRA-based methods, less important ranks still occupy the same parameter budget as important ones. Here, a question is about to arise: Could we free up space from less important ranks for more important ones to achieve better optimization?
@@ 3 BeamLoRA
For a LoRA module with rank $r$, we treat it as a beam with width $r$ and the optimization process is naturally regarded as a search for the solution set $\Delta\mathbf{W}=\{\Delta\mathbf{w}_{1},\Delta\mathbf{w}_{2},...,\Delta\mathbf{w}_{r}\}$ tailored to the fine-tuning dataset, where $i$-th rank in the LoRA module is considered a sub-solution $\Delta\mathbf{w}_{i}$. Formally, the optimization process seeks to minimize the loss function $\mathcal{L}$ over dataset $\mathcal{D}$:
@@ 3.1 Importance Assessment
At the start of the fine-tuning process, LoRA initializes each rank to zero, indicating that their initial importance is equal. Consequently, we initialize all elements in $\mathbf{s}$ with identical values. During the fine-tuning process, $\mathbf{s}$ is consistently normalized using the softmax function, like the logits of tokens in text generation,
@@ 3.2 Pruning and Expansion
For each pruned rank, we copy the parameter values from the corresponding important rank:
@@
Meanwhile, to ensure the stability of the optimization for the expanded ranks, the optimizer states are also copied:
@@
However, directly copying parameters and optimizer states from the original ranks creates a challenge: the lack of symmetry breaking between the expanded and original ranks means the optimization process is essentially trying to synchronously optimize two identical objects (Chen et al., 2016). This makes it difficult to effectively leverage the additional capacity provided by the expanded parameter space. To address this issue, we propose using historical parameters and their corresponding optimizer states to break the symmetry, Eq. 9 and Eq. 10 change to:
@@ 3.3 Dynamic Top-P Threshold
In the previous statement, we fix the number of ranks for each pruning or extension operation to be $K$. This might overlook the actual distribution of parameter importance, potentially leading to the elimination of relatively important parameters due to quantitative constraints. Similar to the sampling process of text generation, we introduce Top-P strategy (Holtzman et al., 2020; Huang et al., 2024) to dynamically determine the operable rank number. Specifically, given the score vector $\mathbf{s}$ and a threshold $p$, we sort $s_{i}$ in descending order, then identify the subset of operable ranks and its size as $K$:
@@
Table 1: Math reasoning and code generation results for LLaMA2-7B and Mistral-7B with $r=64$ for all methods. On Mistral-7B, we compare the baseline methods that perform well on LLaMA. The math reasoning results for Full-FT are derived from MetaMathQA paper (Yu et al., 2024).
@@ 4.1 Experimental Settings
Table 2: Commonsense reasoning results for LLaMA2-7B and LLaMA2-13B. We set $r=32$ for all methods. All results with are taken from Liu et al. (2024) and those marked with are taken from Wang et al. (2024).
@@ 4.2 Math Reasoning and Code Generation
Furthermore, we extend our experiments on Mistral-7B by comparing BeamLoRA with the baseline methods that perform well in LLaMA experiments. The results show that BeamLoRA continues to outperform all baseline methods, surpassing Full-FT by 0.42%. More notably, BeamLoRA shows improvement over Full-FT across all task metrics. This demonstrates that within a limited parameter budget, BeamLoRA can effectively achieve better optimization by expanding the parameter space of important ranks.
@@ 4.3 Commonsense Reasoning
Table 2 presents evaluation results across 8 commonsense reasoning datasets. BeamLoRA achieves the best overall performance on LLaMA2-7B, with an average accuracy improvement of 3.2% over original LoRA and 1.1% over the strong baseline DoRA. Similar to math and coding tasks, BeamLoRA’s performance does not rely heavily on optimal results from just two or three datasets, as observed in other baselines. Instead, it consistently performs among the top three results across most datasets, demonstrating the generalization capability of BeamLoRA. Additionally, we find that IncreLoRA and AdaLoRA are less effective in commonsense reasoning tasks, likely due to frequent rank changes across modules, which cause instability in fine-tuning. This issue is more evident in scenarios requiring extensive task evaluation.
For larger base models LLaMA2-13B, the performance gaps between different methods become smaller. In this setting, BeamLoRA still achieves a 0.4% performance improvement over LoRA, while other methods show inferior performance compared to LoRA. This demonstrates that BeamLoRA can effectively enhance LoRA’s performance with larger base models.
@@ 4.4 Ablation Study
Figure 4: Comparison of BeamLoRA and LoRA on importance of ranks within a module by pruning ranks. The solid lines represent the accuracy changes after pruning, while the dashed lines indicate the accuracy change trends caused by pruning.
@@ 4.5 Analysis
Figure 5: Comparison of BeamLoRA and LoRA on accuracy under various rank settings on LLaMA2-7B.
5 Related Works
@@ 5.1 LoRA and its variants
As one of the parameter-efficient fine-tuning methods, LoRA (Hu et al., 2022) has been widely adopted. However, it still has room for improvement in terms of accuracy. Current enhancements follow two main pathways: optimizing initialization and refining the fine-tuning process. For initialization, methods like PiSSA (Meng et al., 2024) and MiLoRA (Wang et al., 2024) use Singular Value Decomposition on base model weights, with PiSSA focusing on principal singular values and MiLoRA on minor ones for initializing LoRA before fine-tuning. For fine-tuning, DoRA (Liu et al., 2024) splits LoRA’s fine-tuning into magnitude and direction components. ReLoRA (Lialin et al., 2024) continuously merges the fine-tuned LoRA modules into the base model. AdaLoRA (Zhang et al., 2023b) and IncreLoRA (Zhang et al., 2023a) optimize rank allocation across modules. Unlike these approaches, BeamLoRA revisits the foundational aspects of LoRA and recognizes the varying importance of ranks within a module. It compresses less important ranks to free up space for expanding the important ones, thereby allowing them to be better optimized.
@@ 5.2 Model Pruning and Expansion
Model pruning is typically used to remove redundant parameters in models, thereby improving efficiency (Kurtic et al., 2022; Ma et al., 2023). Unlike previous works, our primary goal for pruning is to free up space for expanding important parameters. Model Width expansion is first introduced by Net2Net (Chen et al., 2016) and applied to CNNs. bert2BERT (Chen et al., 2022) extends this method to the pre-training of language models, and the recent work Scaling Smart (Samragh et al., 2024) applies width expansion to large scale base models. Unlike these approaches, we focus on parameter-efficient fine-tuning and propose compressing unimportant parameters within a limited space to expand important ones for better performance. Additionally, due to the shorter nature of the fine-tuning process than pre-training, we propose to use historical states to break symmetry in expansion, thereby ensuring fast convergence.
@@ 6 Conclusion
This paper introduces a PEFT method called BeamLoRA. We adopt a novel perspective to study the characteristics of ranks within a LoRA module and discover that there are differences in the importance of ranks, which change with the number of fine-tuning steps. Based on this, we propose using a dynamic threshold to prune less important ranks, free up space to better optimize the more important ones. Extensive experiments demonstrate that BeamLoRA effectively enhances the performance of LoRA across different base models, tasks, and settings, outperforming other baseline methods.
@@
BeamLoRA introduces a method that compresses less important ranks while expanding important ones during fine-tuning. Although this approach achieves good performance in parameter-efficient fine-tuning, existing implementation requires the addition of a trainable assessment vector $\mathbf{s}$ over ranks. In the context of full-parameter training, a full-rank matrix does not have the low-rank structure like the $\mathbf{B}$-$\mathbf{A}$ decomposition in LoRA, making it impossible to add the vector $\mathbf{s}$. How to extend this method to full-parameter training scenarios remains an area for future research exploration.
@@ A.2 Implementation Details
In the math and code tasks, we follow Yu et al. (2024) and set the Full-FT learning rate for LLaMA2-7B to 2e-5 and for Mistral-7B to 5e-6, with a batch size of 128. The models are fine-tuned for 3 epochs on the dataset. For the ReLoRA method, we use the same settings as Full-FT. For other PEFT methods, we add the LoRA module to all linear layers. Following Wang et al. (2024) we set their learning rate to 3e-4, a batch size of 32, and fine-tuning for 3 epochs.
In the commonsense reasoning tasks, to facilitate comparison, we follow the approach of Hu et al. (2023a) and Liu et al. (2024) by adding LoRA to q_proj, k_proj, v_proj, up_proj, and down_proj modules. We use the optimal learning rate from {2e-4,3e-4} for different methods on LLaMA2-7B, with a batch size of 16 and fine-tuning for 3 epochs. For the larger LLaMA2-13B, following Liu et al. (2024), we reduce the learning rate to the optimal {1e-4,2e-4,3e-4} for different methods, while keeping other settings unchanged. These settings are applied to all PEFT methods. All our experiments are conducted on four H800 GPUs.

### v2

@@ Abstract
Due to the demand for efficient fine-tuning of large language models, Low-Rank Adaptation (LoRA) has been widely adopted as one of the most effective parameter-efficient fine-tuning methods. Nevertheless, while LoRA improves efficiency, there remains room for improvement in accuracy. Herein, we adopt a novel perspective to assess the characteristics of LoRA ranks. The results reveal that different ranks within the LoRA modules not only exhibit varying levels of importance but also evolve dynamically throughout the fine-tuning process, which may limit the performance of LoRA. Based on these findings, we propose BeamLoRA, which conceptualizes each LoRA module as a beam where each rank naturally corresponds to a potential sub-solution, and the fine-tuning process becomes a search for the optimal sub-solution combination. BeamLoRA dynamically eliminates underperforming sub-solutions while expanding the parameter space for promising ones, enhancing performance with a fixed rank. Extensive experiments across three base models and 12 datasets spanning math reasoning, code generation, and commonsense reasoning demonstrate that BeamLoRA consistently enhances the performance of LoRA, surpassing the other baseline methods [^1] The code is available at https://github.com/gccnlp/BeamLoRA..
@@ 1 Introduction
In recent years, large language models have shown tremendous potential in various applications (Touvron et al., 2023a, b; Jiang et al., 2023; OpenAI, 2023; Chen et al., 2024a; Qwen et al., 2025). To further enhance model performance on specific downstream tasks, fine-tuning is usually the most effective approach. However, as the scale of models keeps increasing, fine-tuning all model parameters becomes unsustainable. To address this issue, parameter-efficient fine-tuning (PEFT) emerges as a practical solution (Houlsby et al., 2019; Li and Liang, 2021; Liu et al., 2022; Hu et al., 2022; Gu et al., 2024; Jiang et al., 2024). By updating only lightweight modules, these methods nearly achieve the results of full parameter fine-tuning while reducing both fine-tuning time and memory usage.
@@ 2.1 Low-Rank Adaptation (LoRA)
where $x$ represents the input and $y$ is the output. During fine-tuning, $\mathbf{W}_{0}$ remains frozen, while only $\mathbf{B}$ and $\mathbf{A}$ matrices are trainable.
@@ 2.2 Analysis of LoRA Ranks
Spatial Dimension. Figure 1 describes the results of sorting the importance of ranks after fine-tuning. The blue line represents the levels of importance across different deciles after being sorted by importance in a LoRA module. It can be observed that the deciles of the importance of ranks are hierarchical, indicating that the importance of different ranks within a LoRA module is not uniform after fine-tuning. Furthermore, by pruning the ranks in each LoRA module based on its own importance from the least to the most significant gradually (the red line), the accuracy shows limited change when pruning the less important ranks. On the contrary, when important ranks are pruned, the evaluation results drop sharply to zero. This phenomenon further demonstrates the significant differences in importance among the different ranks.
@@
Figure 1: Differences in importance among ranks within a LoRA module (spatial). The blue line represents the deciles of importance for ranks. The red line represents accuracy changes when pruning ranks of varying importance. We take ffn.up_proj in the 30th layer as an example, with similar phenomena in other modules. By pruning ranks, it can be observed that pruning less important ranks has a minor effect, while pruning more important ranks has a significant impact.
@@
Figure 2: The visualization of importance difference among different ranks in LoRA on LLaMA2-7B as fine-tuning steps increase (temporal). We take ffn.up_proj and attn.v_proj in the 30th layer as examples, with similar trends observed in all other modules. It can be observed that the importance of ranks is equal at the beginning of fine-tuning, but differences gradually emerge as the fine-tuning process progresses.
@@
In summary, there are significant differences in the importance of ranks in LoRA, and these differences appear and gradually increase as the fine-tuning process progresses. However, in most existing LoRA-based methods, less important ranks still occupy the same parameter budget as important ones. Here, a question is about to arise:
Could we free up space from less important ranks for more important ones to achieve better optimization?
@@ 3 BeamLoRA
For a LoRA module with rank $r$, we treat it as a beam with width $r$ and the optimization process is naturally regarded as a search for the solution set $\Delta\mathbf{W}=\{\Delta\mathbf{w}_{1},\Delta\mathbf{w}_{2},...,\Delta\mathbf{w}_{r}\}$ tailored to the fine-tuning dataset, where the $i$-th rank in the LoRA module is considered a sub-solution $\Delta\mathbf{w}_{i}$. Formally, the optimization process seeks to minimize the loss function $\mathcal{L}$ over the dataset $\mathcal{D}$:
@@ 3.1 Importance Assessment
At the start of the fine-tuning process, LoRA initializes each rank to zero, indicating that their initial importance is equal. Consequently, we initialize all elements in $\mathbf{s}$ with identical values. During the fine-tuning process, $\mathbf{s}$ is consistently normalized using the softmax function, like the logits of tokens in text generation:
@@ 3.2 Pruning and Expansion
For the pruned ranks, we assign the parameter values of the important ranks to them:
@@
Meanwhile, to ensure the stability of the optimization for the expanded ranks, the corresponding optimizer states are also assigned:
@@
However, directly assigning parameters and optimizer states from the original ranks creates a challenge: the lack of symmetry breaking between the expanded and original ranks means the optimization process is essentially trying to synchronously optimize two identical objects (Chen et al., 2016). This makes it difficult to effectively leverage the additional capacity provided by the expanded parameter space. To address this issue, we propose using historical parameters and their corresponding optimizer states to break the symmetry, Eq. 9 and Eq. 10 change to:
@@ 3.3 Dynamic Top-P Threshold
In the previous statement, we fix the number of ranks for each pruning or expansion operation to be $K$. This might overlook the actual distribution of parameter importance, potentially leading to the elimination of relatively important parameters due to quantitative constraints. Similar to the sampling process of text generation, we introduce the Top-P strategy (Holtzman et al., 2020; Huang et al., 2024) to dynamically determine the number of operable ranks. Specifically, given the score vector $\mathbf{s}$ and a threshold $p$, we sort $s_{i}$ in descending order, then identify the set of operable ranks and its size as $K$:
@@
Table 1: Math reasoning and code generation results for LLaMA2-7B and Mistral-7B with $r=64$ for all methods. The results are averaged over three runs. On Mistral-7B, we compare the baseline methods that perform well on LLaMA. The math reasoning results for Full-FT are derived from the MetaMathQA paper (Yu et al., 2024).
@@ 4.1 Experimental Settings
Table 2: Commonsense reasoning results for LLaMA2-7B and LLaMA2-13B with $r=32$ for all methods. All results with are taken from DoRA Liu et al. (2024) and those with are taken from MiLoRA Wang et al. (2024).
@@ 4.2 Math Reasoning and Code Generation
Furthermore, we extend our experiments on Mistral-7B by comparing BeamLoRA with the baseline methods that perform well in LLaMA2-7B experiments. The results show that BeamLoRA continues to outperform all baseline methods, surpassing Full-FT by 0.42%. More notably, BeamLoRA shows improvement over Full-FT across all task metrics. This demonstrates that within a limited parameter budget, BeamLoRA can effectively achieve better optimization by expanding the parameter space of important ranks.
@@ 4.3 Commonsense Reasoning
Table 2 presents evaluation results across eight commonsense reasoning datasets. BeamLoRA achieves the best overall performance on LLaMA2-7B, with an average accuracy improvement of 3.2% over original LoRA and 1.1% over the strong baseline DoRA. Similar to math and coding tasks, BeamLoRA’s performance does not rely heavily on optimal results from just two or three datasets, as observed in other baselines. Instead, it consistently performs among the top three results across most datasets, demonstrating the generalization capability of BeamLoRA. Additionally, we find that IncreLoRA and AdaLoRA are less effective in commonsense reasoning tasks, likely due to frequent rank changes across modules, which cause instability in fine-tuning. This issue is more evident in scenarios requiring extensive task evaluation.
For larger base models LLaMA2-13B, the performance gaps between different methods become smaller. In this setting, BeamLoRA still achieves a 0.3% performance improvement over LoRA, while other methods show inferior performance compared to LoRA. This demonstrates that BeamLoRA’s approach of expanding the parameter space of important ranks is effective across base models of different sizes and can enhance the fine-tuning performance of LoRA.
@@ 4.4 Ablation Study
Figure 4: Comparison of BeamLoRA and LoRA on importance of ranks within a module by pruning ranks. The solid lines represent the accuracy changes after pruning, while the dashed lines indicate the accuracy change trends caused by pruning. It can be observed that the importance is more evenly distributed among different ranks in BeamLoRA.
@@ 5 Related Work
Figure 5: Comparison of BeamLoRA and LoRA on accuracy under various rank settings on LLaMA2-7B. BeamLoRA demonstrates consistent improvements over LoRA in all settings.
@@ 5.1 LoRA and Its Variants
As one of the parameter-efficient fine-tuning methods, LoRA (Hu et al., 2022) has been widely adopted. However, it still has room for improvement in terms of accuracy. Recent enhancements follow two main pathways: optimizing initialization and refining the fine-tuning process. For initialization, PiSSA (Meng et al., 2024) and MiLoRA (Wang et al., 2024) apply Singular Value Decomposition on base model weights to initialize LoRA modules, with PiSSA utilizing principal singular values and MiLoRA leveraging minor ones. For fine-tuning, DoRA (Liu et al., 2024) splits LoRA’s fine-tuning into magnitude and direction. ReLoRA (Lialin et al., 2024) continuously merges the fine-tuned LoRA modules into the base model. AdaLoRA (Zhang et al., 2023b) and IncreLoRA (Zhang et al., 2023a) optimize rank allocation across modules. Unlike these approaches, BeamLoRA revisits the foundational aspects of LoRA and recognizes the varying importance of ranks within a module. It compresses less important ranks to free up space for expanding the important ones, allowing them to be better optimized.
@@ 5.2 Model Pruning and Expansion
Model pruning is typically used to remove redundant parameters in models, thereby improving efficiency (Kurtic et al., 2022; Ma et al., 2023; Chen et al., 2024b). Unlike previous works, our primary goal for pruning is to free up space for expanding important parameters. Model width expansion is first introduced by Net2Net (Chen et al., 2016) and applied to CNNs. bert2BERT (Chen et al., 2022) extends this method to the pre-training of language models, and the recent work Scaling Smart (Samragh et al., 2024) applies width expansion to large scale base models. Unlike these approaches, we focus on parameter-efficient fine-tuning and propose compressing unimportant parameters within a limited space to expand important ones for better performance. Additionally, due to the shorter nature of the fine-tuning process than pre-training, we propose to use historical states to break symmetry in expansion, thereby ensuring fast convergence.
@@ 6 Conclusion
This paper introduces a PEFT method called BeamLoRA. We adopt a novel perspective to study the characteristics of ranks within a LoRA module and discover that there are differences in the importance of ranks, which change with the number of fine-tuning steps. Based on this, we propose using a dynamic threshold to prune less important ranks, freeing up space to better optimize the more important ones. Extensive experiments demonstrate that BeamLoRA effectively enhances the performance of LoRA across different base models, tasks, and settings, outperforming other baseline methods.
@@
BeamLoRA introduces a method that compresses less important ranks while expanding important ones during fine-tuning. Although this approach achieves good performance in parameter-efficient fine-tuning, the existing implementation requires the addition of a trainable assessment vector $\mathbf{s}$ over ranks. In the context of full-parameter training, a full-rank matrix does not have the low-rank structure like the $\mathbf{B}$-$\mathbf{A}$ decomposition in LoRA, making it impossible to add the vector $\mathbf{s}$. How to extend this method to full-parameter training scenarios remains an area for future research exploration.
Acknowledgments
The authors thank Yilong Chen from the Institute of Information Engineering for the help. This work is supported by the National Natural Science Foundation of China (No. 62472419, 62472420).
@@ A.2 Implementation Details
In the math and code tasks, we follow Yu et al. (2024) and set the Full-FT learning rate for LLaMA2-7B to 2e-5 and for Mistral-7B to 5e-6, with a batch size of 128. The models are fine-tuned for 3 epochs on the dataset. For the ReLoRA method, we use the same settings as Full-FT. For other PEFT methods, we add the LoRA module to all linear layers. Following Wang et al. (2024) we set their learning rate to 3e-4 on LLaMA2-7B and 6e-5 on Mistral-7B, with a batch size of 32 and fine-tuning for 3 epochs. Our evaluations are conducted using the MetaMathQA and the evalplus Liu et al. (2023) codebase.
In the commonsense reasoning tasks, to facilitate comparison, we follow the approach of Hu et al. (2023a) and Liu et al. (2024) by adding LoRA to q_proj, k_proj, v_proj, up_proj, and down_proj modules. We use the optimal learning rate from {2e-4,3e-4} for different methods on LLaMA2-7B, with a batch size of 16 and fine-tuning for 3 epochs. For the larger LLaMA2-13B, following Liu et al. (2024), we reduce the learning rate to the optimal {1e-4,2e-4,3e-4} for different methods, while keeping other settings unchanged. These settings are applied to all PEFT methods. Our evaluations are conducted using the LLM-Adapters Hu et al. (2023b) codebase. All our experiments are conducted on four H800 GPUs.